Calculate the energy required to ionize a hydrogen atom to an excited state where the electron is initially in the n = 5 energy level. Report your answer in kilojoules.
What is the longest wavelength of electromagnetic radiation capable of ionizing this hydrogen atom in an excited state? Report your answer in nanometers.
1
Expert's answer
2019-09-09T04:59:36-0400

Energy of electron in any energy level n is given by"\frac{-13.6}{n^2}\ eV"

In 5th energy level

"E=\frac{-13.6}{5^2}eV"

Energy required"=E" = "\frac{13.6}{25}=0.544\ eV=1.602\times 10^{-19}\times 0.544\times 10^{-3}\ kJ=8.71\times10^{-23}\ kJ"

Now

"E=\frac{hc}{\lambda}"

"\lambda=\frac{hc}{E}" "=\frac{6.626\times 10^{-34}\times 3\times 10^8}{8.71\times 10^{-23}\times 10^3}=" "2.282\times 10^{-6}\ m=2282\times 10^{-9}m=2282 \ nm"

Maximum Wavelength"=2282 \ nm"